OpenStudy (samigupta8):

integrate the following expression ∫x^3-1/x-1 dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?}(x ^{3}-1)dx/(x-1)=\int\limits_{?}^{?}[(x-1)(x ^{2}+x+1)]dx/(x-1)\] now solve it further..